To resolve an object in an electron microscope, the wavelength of the electrons must be close to the diameter of the object. What kinetic energy, ?k , must the electrons have in order to resolve a protein molecule that is 7.80 nm in diameter? Take the mass of an electron to be 9.11×10−31 kg .
Answer #1

Wavelength (lemda)=7.80 nm =7.80*10^-9 m

De Broglie equation :lemda =h/mv

Speed of electron v=h / m lemda

=6.626*10^-34 / (9.11*10^-31 * 7.80*10^-9 )

= 9.325*10^4 m/s

K.E=1/2 mv^2

m=mass

v=speed

=1/2 *9.11*10^-31*(9.325*10^4)^2

=3.96*10^-21 J ( or 0.247 eV) Answer

I hope help you !!